4b-5=-t for the b
+5 +5
4b=-t+5
4b/4=t+5/4
b= -t+5/4
is this right?

Answer by tutorcecilia(2152)

4b-5=-t
4b-5+5=-t+5
4b=-t+5
4b/4=(-t+5)/4
b=(-t+5)/4
.
check by plugging [b=(-t+5)/4] back into the original equation.
4(-t+5/4)-5=-t
4(-t+5)/4-5=-t
-t+5-5=-t
-t=-t [your answer checks out]